Dwight Hardware and Lumber will allocate the Personnel and Payroll Department's costs to the sales departments of Machines, Automotive, and Furniture. Personnel and Payroll also provide services to each other. Personnel costs are allocated by number of employees, and payroll costs are allocated by gross payroll dollars. Costs and other information for January were as follows:

Personnel Payroll Machines Automotive Furniture
Current costs $13,800 $6,400 $24,400 $40,000 $31,500
Gross Payroll $3,000 $1,500 $5,600 $8,700 $4,050
Number of employees 5 3 8 15 4

Required

a. Prepare a  schedule, which includes the total cost of operating the sales departments for January. Allocate service costs using the step-down method with the sequence of allocation based on highest percentage support concept.
Answer #1

Machine .............30454.28

Automotive ..........50588.61

Furniture ...............35057.11

Explanation:

Highest percentage

Personnel to payroll = 13800 * 3/30 = 1380 ( 3/10 is the ratio based on number of employees )

Payroll to personnel = 6400 * 3000 / 21350 = 899 ( on the basis of wages in dollars )

We see that highest percentage = Personnel department. Hence we allocate the this department cost first and then payroll.

Departments Personnel Payroll Machine Automotive Furniture Current Personnel Payroll Total Cost Allocation Allocation Cost 13